Nuclear forces are mediated by exchange of pi-mesons of REST MASS 140 Mev .
using HEISENBERG UNCERTAINTY principle ,estimate the range OF Nuclear forces.
1
Expert's answer
2020-01-09T10:40:33-0500

Using uncertainty principle in this form


"\Delta E \Delta t\backsim \hbar"

where

"\Delta E=140MeV; \Delta t=\frac {r}{c}"; r - range; c - light velocity

The range of nuclear forces


"r\backsim \frac{\hbar c}{\Delta E }""r\backsim \frac{6.58\cdot 10^{-16}\cdot 3\cdot 10^8}{140\cdot 10^6}=0.141\cdot10^{-14} (m)=1.41\cdot 10^{-13} (cm)"